# Start from the End: A Framework for Computational Policy Exploration to Inform Effective and Geospatially Consistent Interventions applied to COVID-19 in St. Louis

### v1

@@ 2.1 TRACE-STL Model Description
Our complex simulation setting is an updated version of the “TRACE-STL” agent-based model of COVID-19, parameterized and calibrated to the “Omicron wave” of disease spread from December 21, 2021 to February 19, 2022. TRACE-STL was originally developed to forecast robust policy response scenarios over a wide range of potential interventions and epidemiological scenarios [24]. A full description may be found in the Supplement, but we recap the main components and model dynamics here. Briefly, TRACE-STL is a large-scale computational disease transmission model of 2.6 million agents in St. Louis, Missouri and its adjoining counties. We use publicly available data on the number of cases, deaths, and tests [60, 12, 56] , mobility trends [21], vaccinations and boosting [17], and census-tract SVI [6] and simulate a 60-day period where each model timestep denotes one simulated day.
@@ 2.2 Policy Interventions
Vaccines Minimum % of each age group      (5 and older) vaccinated 0 - 75% of each age group
Boosters Minimum % of each age group      (5 and older) boosted 0 - 50% of each age group
@@
Behavior
Symptomatic Testing OR Odds ratio to test symptomatic agents      (relative to non-symptomatic) 10 - 100
Testing Quarantine OR Odds ratio to test quarantined agents      (relative to non-quarantined) 1 - 100
Quarantine Adherence Contact Traced Quarantine adherence when successfully       contact traced 0.7 - 1.0
@@ 3.4 Pooling policies demonstrates a large landscape of small policies to mitigate spread
Figure 4: Pooling mixtures of many low-intensity policies estimated to meet policy goals. The x-axis shows the number of policies “active” at one time (set to a higher value than baseline), and the y-axis shows the percentage of emulated policies estimated to meet the policy goal of 500,000 cumulative infections or less. Each bar is shaded according to the average policy intensity (0 representing baseline, and 1 representing the largest policy value) of the active policies. For k =10 (in the unconstrained case), we emulate 500,000 potential policies. The rightmost bar represents the scenarios where all ten policies were varied, but only consider candidate interventions with no more than 33,000 contact traced per day and mask adherence no more than 0.1 (half of the range explored for each policy). All other bars contain the results of 5,000*$\binom{10}{k}$ policy emulations where k is the number of active policies.
@@ 3.5 Policy specification can reduce disease spread while improving geospatial consistency in outcomes
policy combination $\text{policy combination intensity}=\sum_{i=1}^{10}p_{i}^{2}$
Where $p_{i}$ represents each individual policy strength on a [0-1] scale, and $p_{i}=0\text{ }\forall i$ corresponds to baseline. Note that this differs slightly from the previous analysis where we reported the mean policy intensity: $\frac{1}{k}\sum_{i=1}^{k}p_{i}$ where $p_{i}>0$. This was chosen as a preference for policy combinations that had a mixture of multiple policies (for a two-dimensional example, this is equivalent to saying that combining two policies at half strength would be preferable to doing one at full strength, i.e. that $10.5^{2}+0.5^{2}=0.5\leq 0^{2}+1^{2}=$ 1).

### v2

@@ 2.1 TRACE-STL Model Description
Our complex simulation setting is an updated version of the “TRACE-STL” agent-based model of COVID-19, parameterized and calibrated to the “Omicron wave” of disease spread from December 21, 2021 to February 19, 2022. TRACE-STL was originally developed to forecast robust policy response scenarios over a wide range of potential interventions and epidemiological scenarios [24]. A full description may be found in the Supplement, but we recap the main components and model dynamics here. Briefly, TRACE-STL is a large-scale computational disease transmission model of 2.4 million agents in St. Louis, Missouri and its adjoining counties. The synthetic population and contact networks are based on the 2010 RTI Synthetic population data [65, 32]. We use publicly available data on the number of cases, deaths, and tests [61, 12, 57] , mobility trends [21], vaccinations and boosting [17], and census-tract SVI [6] and simulate a 60-day period where each model timestep denotes one simulated day.
@@ 2.2 Policy Interventions
Vaccines
Minimum % of each age group
   (5 and older) vaccinated
0 - 75% of each age group
Boosters
Minimum % of each age group
   (5 and older) boosted
0 - 50% of each age group
@@
[…]
Symptomatic Testing OR
Odds ratio to test symptomatic agents
   (relative to non-symptomatic)
10 - 100
Testing Quarantine OR
Odds ratio to test quarantined agents
   (relative to non-quarantined)
1 - 100
Quarantine Adherence Contact Traced
Quarantine adherence when successfully
    contact traced
0.7 - 1.0
@@ 3.4 Pooling policies demonstrates a large landscape of small policies to mitigate spread
Figure 4: Pooling mixtures of many low-intensity policies estimated to meet policy goals. The x-axis shows the number of policies “active” at one time (set to a higher value than baseline), and the y-axis shows the percentage of emulated policies estimated to meet the policy goal of 500,000 cumulative infections or less. Each bar is shaded according to the average policy intensity (0 representing baseline, and 1 representing the largest policy value) of the active policies. For k =10 (in the unconstrained case), we emulate 500,000 potential policies. The rightmost bar represents the scenarios where all ten policies were varied, but only consider candidate interventions with no more than 33,000 contact traced per day and mask adherence no more than 0.1 (half of the range explored for each policy). All other bars contain the results of 5,000*(${10}{k}$ ) policy emulations where k is the number of active policies.
@@ 3.5 Policy specification can reduce disease spread while improving geospatial consistency in outcomes
$\text{policy combination intensity}=\sum_{i=1}^{10}p_{i}^{2}$
Where $p_{i}$ represents each individual policy strength on a [0-1] scale, and $p_{i}=0\text{ }\forall i$ corresponds to baseline. Note that this differs slightly from the previous analysis where we reported the mean policy intensity: $\frac{1}{k}\sum_{i=1}^{k}p_{i}$ where $p_{i}>0$. This was chosen as a preference for policy combinations that had a mixture of multiple policies (for a two-dimensional example, this is equivalent to saying that combining two policies at half strength would be preferable to doing one at full strength, i.e. that $0.5^{2}+0.5^{2}=0.5\leq 0^{2}+1^{2}=1$).